OpenStudy (anonymous):

a jar of coins, the ratio of nickels to dimes is 5:7. If the jar contains 350 dimes, how many nickels are there? A) 200 B) 225 C) 250 D) 500

OpenStudy (triciaal):

each nickel is 0.05 each dime is 0.10 350 dimes = 350*0.10 = n/d = 5/7 7/12 = what you just computed for dimes

OpenStudy (triciaal):